OpenStudy (anonymous):

A two-way frequency table shows baldness in men over 45 and men under 45. Under 45 Over 45 Total Bald 24 16 40 Not Bald 36 24 60 60 40 100 Based on this data, are baldness and being over 45 independent events? A) Yes, P(bald | over 45) = P(bald) B) Yes, P(bald | over 45) = P(over 45) C) No, P(bald | over 45) ≠ P(bald) D) No, P(bald | over 45) ≠ P(over 45)

ganeshie8 (ganeshie8):

A and B are independent, if P(A | B) = P(A)

ganeshie8 (ganeshie8):

to knw whether "baldness" and "over 45" are independent or not, find P(bald) first

ganeshie8 (ganeshie8):

p(bald) = ?

OpenStudy (anonymous):

I have no idea what p(bald) is, what does that mean?

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

p(bald) means, probability for being "bald"

ganeshie8 (ganeshie8):

look at the given data, look at last column

ganeshie8 (ganeshie8):

we have surveyed a total of "100" people right ?

OpenStudy (anonymous):

yeah and 40 are bald so is p(bald)=4?

ganeshie8 (ganeshie8):

out of which, 40 are "bald" and 60 are not

OpenStudy (anonymous):

40* not 4

ganeshie8 (ganeshie8):

so, probability for being bald = 40/100 = 4/10 = 2/5

ganeshie8 (ganeshie8):

p(bald) = 2/5

ganeshie8 (ganeshie8):

fine ?

OpenStudy (anonymous):

right

ganeshie8 (ganeshie8):

next u need to find the probability for "bald" for ppl over 45 years

ganeshie8 (ganeshie8):

Look at second column

ganeshie8 (ganeshie8):

40 ppl are over 45 right ?

ganeshie8 (ganeshie8):

out of which, 16 ppl are bald and 24 ppl are not

ganeshie8 (ganeshie8):

so probability of baldness, given that they're over 45 = 16/40 = 2/5

ganeshie8 (ganeshie8):

p(bald | over 45) = 2/5

ganeshie8 (ganeshie8):

so we got the same probability for both :- p(bald) = p(bald | over 45) = 2/5

ganeshie8 (ganeshie8):

so, the two events : baldness and over 45 are independent

ganeshie8 (ganeshie8):

see if that makes some sense...

OpenStudy (anonymous):

I think I understand. Is the answer A then? @ganeshie8

ganeshie8 (ganeshie8):

A is \(\large \color{red}{\checkmark}\)

OpenStudy (anonymous):

Thanks so much! :)

OpenStudy (anonymous):

THis was correct (just took the test)! :) Ty

OpenStudy (anonymous):

You explained how to solve it really well XD